OpenStudy (anonymous):

Can someone please help me answer this problem.. \[\sqrt{x+h} - \sqrt{x} \over h\] = \[1 \over \sqrt{x+h} + \sqrt{x}\]

OpenStudy (anonymous):

\[\sqrt{x+h} - \sqrt{x} \over h\] = \[1 \over \sqrt{x+h} + \sqrt{x}\]

hero (hero):

What do you have to do?

hero (hero):

You can just post something without posting the actual question...

OpenStudy (anonymous):

i have to simplify the equation

OpenStudy (anonymous):

are they supposed to be equal? unlikely

OpenStudy (anonymous):

first one looks like the difference quotient for \[f(x)=\sqrt{x}\]

OpenStudy (anonymous):

like the second equation is supposed to be to the right of the first one..i just didnt know how to type it in

OpenStudy (anonymous):

h = h

OpenStudy (anonymous):

maybe you are supposed to multiply them. that would make more sense

OpenStudy (anonymous):

im supposed to show that they are equivalent

OpenStudy (anonymous):

101ryan101 how did you get h=h?